OpenStudy (anonymous):

I have this question that I quite don't understand so please someone explain to me how to do it! :D Maria began hiking o a trail at a rate of 4 miles per hour for 30 minutes. For the next 1 hour and 15 minutes, she hiked at a rate of 3 miles per hour and completed the trail. What is the total distance Maria hiked?

jimthompson5910 (jim_thompson5910):

"Maria began hiking a trail at a rate of 4 miles per hour for 30 minutes" so how far has she gone?

OpenStudy (anonymous):

2 miles

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

I know that the 1 hour and 15 minutes got me.

jimthompson5910 (jim_thompson5910):

15 min = 15/60 = 1/4 = 0.25 hours So 1 hr, 15 min = 1 hr + 15 min = 1 hr + 0.25 hr = 1.25 hr

jimthompson5910 (jim_thompson5910):

I'm converting to hours because the speed is given in miles per hour

OpenStudy (anonymous):

Ok i dont understand.

jimthompson5910 (jim_thompson5910):

do you see how 15 min is 0.25 hrs?

OpenStudy (anonymous):

How do i convert 1 hour and 15 minutes?

OpenStudy (anonymous):

Yes I see that.

jimthompson5910 (jim_thompson5910):

ok so 1 hr and 15 min is equal to 1 hr + 15 min or 1 hr + 0.25 hr

jimthompson5910 (jim_thompson5910):

Then you add 1 and 0.25 to get 1.25

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Yea i see that.

jimthompson5910 (jim_thompson5910):

ok so 1 hr, 15 min = 1.25 hours

OpenStudy (anonymous):

Oh so in total i add 1.25 and 2?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

where are you getting 2?

jimthompson5910 (jim_thompson5910):

oh nvm i see

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

So i do add them

jimthompson5910 (jim_thompson5910):

no what you do is find the distance she traveled she traveled at 3 mph for 1.25 hrs

OpenStudy (anonymous):

so i divide 1.25 by 3

jimthompson5910 (jim_thompson5910):

how far did she go if she traveled at 3 mph for 1.25 hrs

OpenStudy (anonymous):

Um

OpenStudy (anonymous):

Do i divide that?

jimthompson5910 (jim_thompson5910):

how did you find 2 miles for the first part?

OpenStudy (anonymous):

I divide 4 by 2

jimthompson5910 (jim_thompson5910):

btw the site is being weird, so you may have to be patient

jimthompson5910 (jim_thompson5910):

i see

jimthompson5910 (jim_thompson5910):

well you should have thought of 30 minutes as 1/2 an hour, or 0.5 hours so 4*(1/2) = 2 or 4*(0.5) = 2

jimthompson5910 (jim_thompson5910):

for the second part, 3*(1.25) = 3.75 is the distance she traveled

OpenStudy (anonymous):

lol i did

OpenStudy (anonymous):

So how exactly do I get the answer of 1 hour and 15 mins

jimthompson5910 (jim_thompson5910):

like I did above, you multiply

jimthompson5910 (jim_thompson5910):

since d = r*t

OpenStudy (anonymous):

r?

OpenStudy (anonymous):

and t?

jimthompson5910 (jim_thompson5910):

r is the rate, t is the time

jimthompson5910 (jim_thompson5910):

rate = speed

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so 1.25 times 3

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

equals 3.75

OpenStudy (anonymous):

So i add 3.75 and 2

jimthompson5910 (jim_thompson5910):

this is only possible once you have converted everything to hours

jimthompson5910 (jim_thompson5910):

yes to get 5.75 as the total distance

OpenStudy (anonymous):

oh ok thank you!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

:D